# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v2

@@ IV.2.1 Pure spin-independent quasiperiodic potential
$t_{j}^{\mathrm{eff}}=0,\quad V_{j}^{\mathrm{eff}}=\frac{4t_{0}^{2}}{E-\Delta_{j-1}}+\Delta_{j}.$ (39)
In this case, all eigenstates are localized as shown in Fig. 5(c1). The mechanism behind this can be understood as follows: when $M_{z}=0$ (i.e., no QP Zeeman potential), the Hamiltonian [Eq. (36)] yields two flat bands, and any finite QP onsite potential $M_{z}$ can fully localize all states. Deviating from the high-symmetry line $t_{0}\neq t_{\mathrm{so}}$ introduces next-nearest-neighbor (NNN) hopping, which is given by
@@
where $t_{\pm}=t_{0}\pm t_{\mathrm{so}}$. The nearest-neighbor hopping remains zero, and on-site potential is modified to
$t_{j}^{\mathrm{eff}}=0,\quad V_{j}^{\mathrm{eff}}=\frac{t_{+}^{2}}{E-\Delta_{j-1}}+\frac{t_{-}^{2}}{E-\Delta_{j+1}}+\Delta_{j}.$
The emergence of this NNN hopping begins to destabilize localization, allowing for the appearance of delocalized states. Although the system is no longer analytically solvable for $t_{0}\neq t_{\rm so}$, the possible phases can still be determined. For sufficiently large (weak) $M_{z}$, the system always enters the localized (extended) phase. On the other hand, for intermediate $M_{z}$ values, the chiral symmetry is explicitly broken by spin-independent QP modulation $M_{z}V_{j}^{\mathrm{d}}\sigma_{0}$, leading to the emergence of mobility edges (MEs) between the extended and localized states. This accounts for extended, localized and L+E phases in Fig. 5(b) at $\eta=0$, which extend to small $\eta$ regime with $\eta<\eta_{c}\approx 0.43$.
@@ IV.2.2 Pure spin-dependent quasiperiodic potential
$t_{j}^{\mathrm{eff}}=\frac{-2t_{0}\Delta_{j}}{E},\quad V_{j}^{\mathrm{eff}}=\frac{4t_{0}^{2}+\Delta_{j}^{2}}{E}.$ (41)
@@
$\gamma=\max\Big{\{}\frac{1}{2}\ln|\frac{M_{z}}{4t_{0}}\Big{|},0\Big{\}}.$ (42)
@@
We now extend this analysis to the case where $t_{0}\neq t_{\mathrm{so}}$, which modifies the effective nearest-neighbor hopping and on-site potential due to the imbalance between $t_{0}$ and $t_{\mathrm{so}}$
$t_{j}^{\mathrm{eff}}=\frac{t_{-}\Delta_{j+1}-t_{+}\Delta_{j}}{E},\quad V_{j}^{\mathrm{eff}}=\frac{2t_{0}^{2}+2t_{\mathrm{so}}^{2}+\Delta_{j}^{2}}{E},$
and more importantly, the imbalance introduces an effective next-nearest-neighbor hopping
@@
The system exhibits the chiral symmetry at $\eta=1$, implying the absence of MEs and the presence of only three pure phases: localized, extended, and critical. The global theory does not strictly apply here, however, because the system has no mobility edges, the phase boundaries can be determined by analyzing the typical eigenstates. For example, by considering the zero-energy states $E=0$, the eigen-equation Eq. (36) becomes
$-t_{0}(\psi_{j+1}+\psi_{j-1})+t_{\mathrm{so}}(-\psi_{j+1}+\psi_{j-1})+\Delta_{j}\psi_{j}=0,$
$-t_{0}(\varphi_{j+1}+\varphi_{j-1})+t_{\mathrm{so}}(\varphi_{j+1}-\varphi_{j-1})+\Delta_{j}\varphi_{j}=0.$
Here $\psi_{j}$ and $\varphi_{j}$ are the wavefunction of the spin-up and spin-down states, and the eigen-equation is expressed as the form of transfer matrix $\big{(}\begin{array}[]{cc}\psi_{j+1},&\psi_{j}\end{array}\big{)}^{\intercal}=A_{j}\big{(}\begin{array}[]{cc}\psi_{j},&\psi_{j-1}\end{array}\big{)}^{\intercal}$ and $\big{(}\begin{array}[]{cc}\varphi_{j+1},&\varphi_{j}\end{array}\big{)}^{\intercal}=B_{j}\big{(}\begin{array}[]{cc}\varphi_{j},&\varphi_{j-1}\end{array}\big{)}^{\intercal}$ where the transfer matrix $A_{j}$ is
$A_{j}=\left(\begin{array}[]{cc}\Delta_{j}/{t_{+}}&-t_{-}/t_{+}\\ 1&0\end{array}\right),$
and $B_{j}$ has a similar expression. In the spin-dependent limit ($\eta=1$), the delocalized phase is topological while the localized phase is topologically trivial Wang et al. (2020b). The phase boundary between the critical and localized phases can thus be determined by the topological transition. If both eigenvalues of the transfer matrix $A=\prod_{j=1}^{L}A_{j}$ are either less than 1 or greater than 1, the system is topological, and the $\psi$ zero mode is localized at one end of the chain. Assuming $t_{0}>0$ and $t_{\mathrm{so}}>0$, then the two eigenvalues of $A$ satisfy $|\lambda_{1}\lambda_{2}|<1$ , the topological nature is dictated by the larger eigenvalue $|\lambda_{2}|$. After performing a standard similarity transformation DeGottardi et al. (2011, 2013) or applying the global theory, we find the localized-to-critical transition point is given by $|M_{z}/2t_{+}|=1$. By applying the same reasoning to the transfer matrix in the dual space, the phase boundary between localized and critical phases in the dual space is $|M_{z}/2t_{-}|=1$, corresponding to the transition between extended and critical phases in the original space. Therefore, for $\eta=1$ and $t_{0}\neq t_{\mathrm{so}}$, the system exhibits three distinct phases as summarized in Table. 2.
@@ IV.2.3 Balanced quasiperiodic potential
$t_{j}^{\mathrm{eff}}=\frac{-2t_{0}\Delta_{j-1}}{2E-\Delta_{j-1}},\quad V_{j}^{\mathrm{eff}}=\frac{8t_{0}^{2}}{2E-\Delta_{j-1}}+\frac{E\Delta_{j}}{2E-\Delta_{j}}.$ (44)
@@
Figure 6: Experimental realization schemes. (a)-(b) Implementation of the model supporting all fundamental mobility edges (MEs) (a) The spin-conserved hopping for spin-down atoms is suppressed by a spin-dependent deep primary potential ${\cal V}_{p}(z)$ (blue solid line), while spin-up atoms undergo hopping in a secondary potential ${\cal V}_{s}(z)$ (red solid line). The on-site potential for spin-down atoms (red dashed line) is realized by the combination of ${\cal V}_{p}(z)$ and ${\cal V}_{s}(z)$ , which are incommensurate with each other. The on-site spin-flipped process is introduced via an overall coupling $M_{0}$. (b) Schematic of the experimental setup. The spin-dependent potential ${\cal V}_{p}(z)$ is formed by the combination of ${\bf E}_{2}$ and ${\bf E}_{3}$ with the traveling wave ${\bf E}_{4}$, while the spin-independent secondary potential ${\cal V}_{s}(z)$ is generated by ${\bf E}_{1}$, affecting both spin states. (c)-(d) Realization of the quasiperiodic optical Raman lattice model. (c) The spin-conserved hopping $t_{0}$ is induced by the primary lattice ${\cal V}_{p}(z)$, and the spin-flipped hopping $t_{\rm so}$ is induced by the Raman potential ${\cal M}(z)$. The quasiperiodic Zeeman potential is realized by the secondary lattice ${\cal V}_{s}(z)$. (d) Schematic of the setup. A standing wave ${\bf E}_{1}$, combined with the blue-detuned spin-independent primary lattice generated by ${\bf E}_{2}$, creates the spin-dependent Zeeman potential. A traveling wave ${\bf E}_{3}$ generates the Raman potential ${\cal M}(z)$ in combination with ${\bf E}_{2}$.
@@ IV.3 Unified picture for the seven fundamental localization phases
It is important to highlight the mapping relation between two analytically solvable models: the one used in our experiment and the type-II mosaic lattice model Zhou et al. (2023) defined in Eq. (21), with slightly modified on-site potential, given by
@@
Here, $V_{0}$ is the strength of the on-site potential. At the high symmetry line $V_{0}=t$, the model Eq. (47) has analytic expression for the MEs as discussed in Sec. III.4. Its MEs between critical and localized states, and this mapping provides a systematic framework for understanding the entire phase diagram.
The QP optical lattice model model at $t_{0}=t_{\mathrm{so}}$ can be mapped to the type-II mosaic lattice model [Eq. (47)] followed by a local rotation $U=\exp(-i\pi\sigma_{y}/4)$. The coefficients between the two models are related by the following substitution rule
@@
In the case of a pure spin-dependent QP potential ($\eta=1$) of QP optical Raman lattice model, the type-II mosaic model reduces to the uniform on-site potential limit $V_{0}=0$. In this scenario, both models are exactly solvable, exhibiting localized and critical phases. Specifically, the type-II mosaic model is in the localized phase when $|t|>|\lambda|$ and in the critical phase when $|t|<|\lambda|$ [Fig. 3(a)]. This is consistent with the QP optical Raman lattice model, where the localized phase occurs when $M_{z}>4t_{0}$ and the critical phase when $M_{z}<4t_{0}$ [Fig. 5(c3)].
For the balanced quasiperiodic potential ($\eta=1/2$), the type-II mosaic model is along the high-symmetry line $V_{0}=t$. In this case, both models are exactly solvable again and exhibits only L+C phase, which corresponds to the coexistence of localized and critical states due to the presence of mobility edges. The type-II mosaic model exhibits exact MEs at $E=\pm\lambda$ separating the localized and critical states Zhou et al. (2023), which corresponds to the MEs which are given by $E=\pm 2t_{0}$ in the QP optical Raman lattice model.
Reducing the chiral parameter $\eta$ from 1 to 1/2 corresponds to turning on the on-site potential $V_{0}$ in the type-II mosaic model, transitioning from $V_{0}=0$ to $V_{0}=t$. Breaking the chiral symmetry of both models, introducing the MEs between localized and critical states, leading to the emergence of L+C phase. As the on-site potential increases towards the high-symmetry line $V_{0}=t$, or when the QP optical Raman lattice model reaches $\eta=1/2$, the system becomes analytically solvable again, exhibiting L+C phase in the phase diagram and indicating the presence of MEs between localized and critical states.
Further introducing the imbalance between $t_{0}$ and $t_{\mathrm{so}}$ leads to the emergence of extended phases and associated MEs. And this will lead to the long-range hopping terms when reduced the system into the spinless QP chain, introducing extended orbital into the spectrum of the type-II mosaic model. Thus, this mapping, combined with the imbalance between $t_{0}$ and $t_{\mathrm{so}}$, accounts for the generation of the seven phases in the phase diagram.
@@ V Schemes for experimental realization
In this section, we propose experimental schemes for realizing models that host all three fundamental types of mobility edges (MEs) [Eq.(32)] and all seven fundamental phases [Eq.(36)]. These schemes utilize ultracold alkali atoms, which are ideal candidates due to their large fine-structure energy splitting and moderate natural linewidth. These properties facilitate efficient Raman coupling and spin-dependent potentials, even when the laser frequency is far detuned from the $D_{1}$ and $D_{2}$ lines. Such features have been extensively studied both theoretically and experimentally Soltan-Panahi et al. (2011); Kuzmenko et al. (2019); Szulim et al. (2022), making ultracold alkali atoms well-suited for implementing an incommensurate lattice schemes.
We first propose the scheme for the model in Eq. (27) with $\mu=\lambda_{1}=0$ supporting all three fundamental types of MEs, as illustrated in Fig. 6(a-b), whose Hamiltonian is given by
@@
where the first term represents the kinetic energy. Here ${\cal V}_{p}(z)$ is a deep spin-dependent primary lattice that freezes the spin-conserved hopping for spin-down atoms only, and is given by ${\cal V}_{p}(z)=V_{p}\cos(k_{p}z+\phi_{p})$. The terms in Eq. (32) can be realized as follows. The uniform spin-conserved hopping term $\Pi_{j}=t\Lambda_{+}$ for spin-up atoms can be obtained by
@@
The Hamiltonian in Eq. (51) can be realized based on QP optical Raman lattice, as outlined below. The detailed implementation can be found in Appendix. E. To facilitate the description, we perform a unitary transformation, such that $\sigma_{z}\rightarrow\sigma_{x}$ and $\sigma_{x}\rightarrow-\sigma_{z}$. The kinetic term and spin-independent potential ${\cal V}_{s}(z)$ remain unchanged and can be generated using the standing wave ${\bf E}_{1}$. The rotated spin-dependent potential ${\cal V}_{p}(z)(\sigma_{0}-\sigma_{x})/2$ can be realized using a standing waves ${\bf E}_{2}$ and ${\bf E}_{3}$ along with a traveling wave ${\bf E}_{4}$. Finally, the $M_{0}\sigma_{z}$ term arises from the two-photon detuning of the Raman coupling. After performing the inverse unitary trnasformation, we reach the Hamiltonian in Eq. (51). More details can be found in Appendix. E, where ${}^{87}{\rm Rb}$ atoms are used as an exmaple.

### v4

@@ IV.2.1 Pure spin-independent quasiperiodic potential
$t_{j}^{\mathrm{eff}}=0,\ V_{j}^{\mathrm{eff}}=\frac{4t_{0}^{2}}{E-\Delta_{j-1}}+\Delta_{j}.$ (41)
In this case, all eigenstates are localized as shown in Fig. 5(c1). The mechanism behind this can be understood as follows: when $M_{z}=0$ (i.e., no QP potential), the Hamiltonian [Eq. (38)] yields two flat bands, and any finite QP onsite potential $M_{z}$ can fully localize all states. Deviating from the high-symmetry line $t_{0}\neq t_{\mathrm{so}}$ introduces next-nearest-neighbor (NNN) hopping $t_{j}^{\mathrm{NN}}a_{j}^{\dagger}a_{j+2}+{\rm h.c.}$, with coefficient
@@
where $t_{\pm}=t_{0}\pm t_{\mathrm{so}}$. In this regime, nearest-neighbor hopping remains zero $t_{j}^{\rm eff}=0$, and on-site potential becomes
$V_{j}^{\mathrm{eff}}=\frac{t_{+}^{2}}{E-\Delta_{j-1}}+\frac{t_{-}^{2}}{E-\Delta_{j+1}}+\Delta_{j}.$ (43)
The emergence of NNN hopping begins to destabilize localization, allowing for the appearance of delocalized states. Although the system is no longer analytically solvable for $t_{0}\neq t_{\rm so}$, the possible phases can still be determined. For sufficiently large (weak) $M_{z}$, the system always enters the localized (extended) phase. On the other hand, for intermediate $M_{z}$ values, the chiral symmetry is explicitly broken by spin-independent QP modulation $M_{z}V_{j}^{\mathrm{d}}\sigma_{0}$, leading to the emergence of mobility edges (MEs) between the extended and localized states. This accounts for extended, localized and L+E phases in Fig. 5(b) at $\eta=0$, which extend to small $\eta$ regime with $\eta<\eta_{c}\approx 0.43$.
@@ IV.2.2 Pure spin-dependent quasiperiodic potential
$t_{j}^{\mathrm{eff}}=\frac{-2t_{0}\Delta_{j}}{E},\ V_{j}^{\mathrm{eff}}=\frac{4t_{0}^{2}+\Delta_{j}^{2}}{E}.$ (44)
@@
$\gamma=\max\left\{\frac{1}{2}\ln\left|\frac{M_{z}}{4t_{0}}\right|,0\right\}.$ (45)
@@
We extend our analysis to the case where $t_{0}\neq t_{\mathrm{so}}$, which introduces modifications to the effective nearest-neighbor hopping and on-site potential due to the imbalance between $t_{0}$ and $t_{\mathrm{so}}$
$t_{j}^{\mathrm{eff}}=\frac{t_{-}\Delta_{j+1}-t_{+}\Delta_{j}}{E},\ V_{j}^{\mathrm{eff}}=\frac{2t_{0}^{2}+2t_{\mathrm{so}}^{2}+\Delta_{j}^{2}}{E},$ (46)
and more importantly, such imbalance introduces an effective NNN hopping, with coefficient given by
@@
The system exhibits the chiral symmetry at $\eta=1$, leading to the vanishing MEs and the presence of only three pure phases: localized, extended, and critical. The global theory does not strictly apply here, however, since the system has no mobility edges, the phase boundaries can be determined by analyzing the typical eigenstates. For example, by considering the zero-energy states $E=0$, the eigen-equation Eq. (38) for spin-up (spin-down) states $\psi_{j}$ ($\varphi_{j}$) becomes
$-t_{0}(\psi_{j+1}+\psi_{j-1})-t_{\mathrm{so}}(\psi_{j+1}-\psi_{j-1})+\Delta_{j}\psi_{j}=0,$ (48)
with the $\varphi$ obtained by replacing $t_{\rm so}\to-t_{\rm so}$. The above equation can be expressed in the form of transfer matrix as $(\psi_{j+1},\psi_{j})^{\intercal}=A_{j}(\psi_{j},\psi_{j-1})^{\intercal}$, with $A_{j}$ being
$A_{j}=\left(\begin{array}[]{cc}\Delta_{j}/{t_{+}}&-t_{-}/t_{+}\\ 1&0\end{array}\right),$ (49)
with an analogous form for spin-down states. In the pure spin-dependent limit ($\eta=1$), the delocalized phase is topological while the localized phase is topologically trivial Wang et al. (2020b). The phase boundary between the critical and localized phases can thus be determined by the topological transition. If both eigenvalues of the transfer matrix $A=\prod_{j=1}^{L}A_{j}$ are either less than 1 or greater than 1, the system is topological, and the $\psi$ zero mode is localized at one end of the chain. Assuming $t_{0}>0$ and $t_{\mathrm{so}}>0$, then the two eigenvalues of $A$ satisfy $|\lambda_{1}\lambda_{2}|<1$ , the topological nature is dictated by the larger eigenvalue $|\lambda_{2}|$. After performing a standard similarity transformation DeGottardi et al. (2011, 2013) or applying the global theory, we find the localized-to-critical transition point is given by $|M_{z}/2t_{+}|=1$. By applying the same reasoning to the transfer matrix in the dual space, the phase boundary between localized and critical phases in the dual space is $|M_{z}/2t_{-}|=1$, corresponding to the transition between extended and critical phases in the original space. Therefore, for $\eta=1$ and $t_{0}\neq t_{\mathrm{so}}$, the system exhibits three distinct pure phases as summarized in Table. 2.
@@ IV.2.3 Balanced quasiperiodic potential
$t_{j}^{\mathrm{eff}}=\frac{-2t_{0}\Delta_{j-1}}{2E-\Delta_{j-1}},\ V_{j}^{\mathrm{eff}}=\frac{8t_{0}^{2}}{2E-\Delta_{j-1}}+\frac{E\Delta_{j}}{2E-\Delta_{j}}.$ (50)
@@
Figure 6: Experimental realization schemes. (a)-(b) Implementation of the model supporting all fundamental mobility edges (MEs). (a) The spin-conserved hopping for spin-down atoms is suppressed by a spin-dependent deep primary potential ${\cal V}_{p}(z)$ (blue solid line), while spin-up atoms undergo hopping in a secondary potential ${\cal V}_{s}(z)$ (red solid line). The on-site potential for spin-down atoms (red dashed line) is realized by the combination of ${\cal V}_{p}(z)$ and ${\cal V}_{s}(z)$ , which are incommensurate with each other. The on-site spin-flipped process is introduced via an overall coupling $M_{0}$. (b) Schematic of the experimental setup. The spin-dependent potential ${\cal V}_{p}(z)$ is formed by the combination of ${\bf E}_{2}$ and ${\bf E}_{3}$ with the traveling wave ${\bf E}_{4}$, while the spin-independent secondary potential ${\cal V}_{s}(z)$ is generated by ${\bf E}_{1}$, affecting both spin states. (c)-(d) Realization of the quasiperiodic optical Raman lattice model. (c) The spin-conserved hopping $t_{0}$ is induced by the primary lattice ${\cal V}_{p}(z)$, and the spin-flipped hopping $t_{\rm so}$ is induced by the Raman potential ${\cal M}(z)$. The quasiperiodic Zeeman potential is realized by the secondary lattice ${\cal V}_{s}(z)$. (d) Schematic of the setup. A standing wave ${\bf E}_{1}$, combined with the blue-detuned spin-independent primary lattice generated by ${\bf E}_{2}$, creates the spin-dependent Zeeman potential. A traveling wave ${\bf E}_{3}$ generates the Raman potential ${\cal M}(z)$ in combination with ${\bf E}_{2}$.
@@ IV.3 Unified picture for the seven fundamental localization phases
We highlight the mapping relation between the present 1D QP optical Raman lattice model in Eq. (38) and the type-II QP mosaic model Zhou et al. (2023) in Eq. (23), with slightly modified on-site potential, given by
@@
Here, $V_{0}$ is the strength of the on-site potential. At the high symmetry line $V_{0}=t$, the model Eq. (53) reduces to the origin one in Eq. (38), with analytic MEs between critical and localized states as discussed in Sec. III.4. This mapping deepens our understanding for the entire phase diagram in Fig. 5(b).
The QP optical lattice model model at $t_{0}=t_{\mathrm{so}}$ can be mapped to the type-II QP mosaic model [Eq. (53)] followed by a local rotation $U=\exp(-i\pi\sigma_{y}/4)$, with the coefficients between two models are related by the following substitution rule
@@
The pure spin-dependent QP potential limit ($\eta=1$) of QP optical Raman lattice model corresponds to the type-II QP mosaic model in the limit $V_{0}=0$. In this scenario, both models are exactly solvable, exhibiting localized and critical phases. Specifically, the type-II QP mosaic model is in the localized phase when $|t|>|\lambda|$ and in the critical phase when $|t|<|\lambda|$ [Fig. 3(a)]. This is consistent with the QP optical Raman lattice model, where the localized phase occurs when $M_{z}>4t_{0}$ and the critical phase when $M_{z}<4t_{0}$ [Fig. 5(c3)].
Reducing the chiral parameter $\eta$ from 1 to 1/2 corresponds to turning on the on-site potential $V_{0}$ in the type-II QP mosaic model, transitioning from $V_{0}=0$ to $V_{0}=t$. Breaking the chiral symmetry of both models, introducing the MEs between localized and critical states, leading to the L+C phase. As the on-site potential increases towards the high-symmetry line $V_{0}=t$, or when the QP optical Raman lattice model reaches $\eta=1/2$, the system becomes analytically solvable again. The type-II QP mosaic model exhibits analytic MEs $E=\pm\lambda$ separating the localized and critical states Zhou et al. (2023), which corresponds to the MEs $E=\pm 2t_{0}$ in the QP optical Raman lattice model.
Further introducing the imbalance between $t_{0}$ and $t_{\mathrm{so}}$ leads to the emergence of extended phases and associated MEs. And this will lead to the long-range hopping terms when reduced the system into the spinless QP chain, introducing extended orbital into the spectrum of the type-II QP mosaic model. Thus, this mapping, combined with the imbalance between $t_{0}$ and $t_{\mathrm{so}}$, accounts for the generation of the seven phases in the phase diagram.
@@ V Schemes for experimental realization
Finally, we propose experimental schemes for realizing the SSQP on-site model [Eq. (34)] and QP optical Raman lattice model [Eq. (38)]. These schemes utilize ultracold alkali atoms, which are ideal candidates due to their large fine-structure energy splitting and moderate natural linewidth. These properties facilitate efficient Raman coupling and spin-dependent potentials, even when the laser frequency is far detuned from the $D_{1}$ and $D_{2}$ lines Soltan-Panahi et al. (2011); Kuzmenko et al. (2019); Szulim et al. (2022), making ultracold alkali atoms well-suited for implementing the QP lattice.
We first propose the scheme for the SSQP on-site model in Eq. (34) supporting all three basic MEs, as illustrated in Fig. 6(a-b), whose Hamiltonian is
@@
where the first term represents the kinetic energy. ${\cal V}_{p}(z)=V_{p}\cos(k_{p}z+\phi_{p})$ is a deep spin-dependent primary lattice that freezes the spin-conserved hopping for spin-down atoms. The different terms in Eq. (34) can be realized as follows. The uniform spin-conserved hopping $\Pi_{j}=t\Lambda_{+}$ for spin-up atoms can be obtained by
@@
The proposed Hamiltonian in Eq. (57) can be realized based on QP optical Raman lattice, as outlined below. The detailed implementation can be found in Appendix. G. To facilitate the description, we perform a unitary transformation, such that $\sigma_{z}\rightarrow\sigma_{x}$ and $\sigma_{x}\rightarrow-\sigma_{z}$. The kinetic term and spin-independent potential ${\cal V}_{s}(z)$ remain unchanged and can be generated using the standing wave ${\bf E}_{1}$. The rotated spin-dependent potential ${\cal V}_{p}(z)(\sigma_{0}-\sigma_{x})/2$ can be realized using a standing waves ${\bf E}_{2}$ and ${\bf E}_{3}$ along with a traveling wave ${\bf E}_{4}$. Finally, the $M_{0}\sigma_{z}$ term arises from the two-photon detuning of the Raman coupling. After performing the inverse unitary transformation, we reach the Hamiltonian in Eq. (57). More details can be found in Appendix. G, where ${}^{87}{\rm Rb}$ atoms are used as an example.